5. Solve the following IVP by using the method of Laplace transforms: w- 2 w +w = 6 t - 2, w(-1) = 3 , w ( -1) = 7
Answer #1

w wz bt- 2 Rebae tb w wwtw l w0) +3, wlo)= Taking labla transfor m Let Llu) ) LwS3 3 5-7 (Q in D Aut w 25)-35-1+6= 6-2s S&9 5where A st6 Yt58- ASCS-5 s-i)+ D 3く45 タ+58 Pot S=1 Dニ 3+1-8+6-2 ut 2 2 AB ) D(4). ニQ4+ 4-16t fut 33A (4) + 8 4)(2)+D()=8r9-a4